OpenStudy (pokemon23):

well technically I don't need math help but I need to see if I have correct answers anyone wanna help me it about factoring...

OpenStudy (anonymous):

go man

OpenStudy (barrycarter):

You can always use wolframalpha.com to check your answers.

OpenStudy (pokemon23):

4x^2-4=4(x^2-1) =4(x-1) (x+1)

OpenStudy (pokemon23):

oh ok

OpenStudy (pokemon23):

so here is my second question

OpenStudy (pokemon23):

which are?